OpenStudy (anonymous):

find the solution to the quadratic equation x^2 + x = 3x - 5

OpenStudy (whpalmer4):

it's right there in front of you! did you mean find the solutions to the quadratic equation?

OpenStudy (anonymous):

oops , yes the solution! find the solution!

OpenStudy (anonymous):

you need to use the quadratic formula on: \[ax^2 + bx + c = 0\] \[x_{1,2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] you get 2 x values. one for plus, one for minus on the plusminus sign

OpenStudy (whpalmer4):

This is also pretty easy to solve by completing the square, if you prefer.

OpenStudy (anonymous):

these are the possible solution A. 5±√21 _______ 2 B. 3±√34 _______ 3 C.-2±√-16 _______ 2 D.2±√-16 _______ 2 E.-2±√21 _______ 4 F.-5±√6 _______ 2

OpenStudy (anonymous):

someone clarify pleeeeasee

OpenStudy (luigi0210):

move everything to one side and then plug in values to the quadratic formula

OpenStudy (anonymous):

my answer is D

OpenStudy (whpalmer4):

D is correct

OpenStudy (dan815):

luigi!

OpenStudy (dan815):

nice score! :)

OpenStudy (jhannybean):

Using @whpalmer4's way, let's try completing the square. \[\large x^2 +x =3x -5\]\[\large x^2 + x - 3x = -5\]\[\large x^2 -2x = -5\]\[ \quad \large c=(\frac{-2}{2})^2=1\]\[\large x^2 - 2x +1 = -5+1\]\[\large (x-1)^2 = -4\]\[\large \sqrt{(x-1)^2}=\sqrt{-4}\]\[\large x-1 = \pm 2i\]\[\large x= 1 \pm2i \implies x- 2i, x+2i\]

OpenStudy (whpalmer4):

I would have stopped at \(x=1\pm2i\), not sure what the rightarrow is supposed to indicate...

OpenStudy (luigi0210):

you will all surpass me eventually, dan :P

OpenStudy (jhannybean):

"implies"

OpenStudy (dan815):

no i wont i will wait for you level then i will lvl

OpenStudy (whpalmer4):

okay, but I don't see how it implies what you have to the right of the arrow :-)

OpenStudy (jhannybean):

@dan815 mine is better. ;) lolol.

OpenStudy (dan815):

69 is that akward phase no one wants to be in

OpenStudy (dan815):

its like being 18 before u can drink

OpenStudy (luigi0210):

well you're gonna have to wait awhile then :P

OpenStudy (jhannybean):

what i mean is if you break it down "x pm 2i" means x=1+ 2i and x= 1-2i lol.

OpenStudy (kenljw):

x^2 + x = 3x - 5 first solve so as to equal zero x^2 -2x +5 = 0 The quadratic takes different forms, in engineering its 0 = x^2 +bx +c = -(b/2)(1 +- sqrt(1 - c(2/b)^2) :b =B/A. c = C/A therefore x = 1(1 +- sqrt(1 - 5(1^2)) = 1 +- i4 when complex always come in conjugate form (x - 1 + i4)(x -1 - i4) = 0

OpenStudy (luigi0210):

@dan815 watch, Jhanny will even pass me.. then she'll be your competition :P

OpenStudy (jhannybean):

I don't feel smart enough!

OpenStudy (whpalmer4):

@kenljw but what about that pesky detail that \[(x-1+4i)(x-1-4i) = x^2-2x+17\]

OpenStudy (whpalmer4):

sqrt(1-5(1^2) = 2i, not 4i

OpenStudy (kenljw):

(1 - i4)(1 +i4) = 1 +i4 -i4 - I^4 = 5

OpenStudy (luigi0210):

Pfft, what are you talking about Jhan, you're beyond smart

OpenStudy (kenljw):

i^2 = -1

OpenStudy (kenljw):

The solution is there, they put in parabola form for drawing. That's the problem when using the computer before you can do by hand. The computer only facilitates the work used be done by hand, slide rule, and calculators.

OpenStudy (kenljw):

I'm old school, I'm 62 EE and gone through them all

OpenStudy (whpalmer4):

@kenljw your solution is incorrect. you forgot to take a square root. don't believe me, try substituting your solution into the original equation.

OpenStudy (whpalmer4):

specifically, this line is incorrect: x = 1(1 +- sqrt(1 - 5(1^2)) = 1 +- i4

OpenStudy (kenljw):

Your correct, that's why always show work so either you can check your arithmetic or someone else